10) Two angles are complementary if their sum is 90e. If the measure of the first angle is xe, and the measure of the
second angle is (3x - 2)e, find the measure of each angle.

Answer by bucky(2189)

You are told that one angle is x degrees and its complement is 3x - 2 degrees.
.
By definition they must add together to total 90 degrees. So you can write the equation:
.
x + 3x - 2 = 90
.
Add the two terms containing x and the equation becomes:
.
4x - 2 = 90
.
Get rid of the -2 on the left side by adding +2 to both sides. The addition of +2 on the
left side cancels the -2 and you get:
.
4x = 92
.
Solve for x by dividing both sides by 4:
.
x = 92/4 = 23 degrees
.
If this angle is 23 degrees, then its complement is 67 degrees because the sum of the two
angles must be 90 degrees.
.
Hope this helps you to understand the problem.
.